In an atom of H, electron is moving in 5th excited state .then, How many possibilities are to get a Balmer line ?  plz explain the concept of excited state in this question....

Grade:12

2 Answers

Vaishal Shah
12 years ago
5th excited state means n=6.....then possible balmer lines are 4..bcoz for balmer n=2.so,possible lines=6-2=4....
basit ali
12 years ago

Hello, Vaishal the answer was correct but i can't understand the concept behind (6-2=4 ) that u had used formula . Can u explain this ? plz explain...

    THANKS!!!!Smile